# ZeroTuning: Unlocking the Initial Token’s Power to Enhance Large Language Models Without Training

### v2

@@ 4.2 Overall performance of ZeroTuning
For a fair and direct comparison with existing supervised baselines, our main experiments focus on the supervised ZeroTuning approach.
@@ 4.3 Unsupervised ZeroTuning
We validate the fully unsupervised variant of ZeroTuning, which eliminates the need for any labeled calibration data by minimizing the model’s average output entropy. Appendix E provides a detailed empirical validation for this approach, including visual comparisons, empirical analysis, and a breakdown of error patterns. As shown in Figure 8, this entropy-guided method achieves performance highly competitive with its supervised counterpart, extending its applicability to label-scarce scenarios.
@@ 5 Further Analysis and Key Findings
Finally, we analyze the method’s boundaries and potential side effects. We quantitatively demonstrate that ZeroTuning excels at correcting a model’s uncertain errors but cannot override high-confidence mistakes rooted in flawed pretrained knowledge. We also characterize the negative effects of extreme tuning, which provides a clearer picture of the method’s operational limits (Appendix D). This positions our method as a powerful tool for unlocking a model’s latent knowledge, rather than a substitute for fine-tuning. Intriguingly, we also find that within a safe operational range, the scaling factor can modulate output diversity in a manner analogous to temperature, but with the unique ability to alter the rank-ordering of logits and thereby correct errors that temperature scaling cannot fix.
@@ 6 Conclusion
In this work, we present a systematic analysis of tuning the initial token’s attention and propose ZeroTuning, a novel, training-free method to enhance LLMs. By recalibrating this single, task-agnostic token, ZeroTuning outperforms previous methods that require task-specific tuning. It operates effectively in both supervised and unsupervised modes and demonstrates broad compatibility across various implementations. This work advances inference-time tuning and contributes to the interpretability of LLMs, opening new avenues for lightweight model optimization.
@@ A.1 Token-Level Attention Tuning
Token-level attention tuning typically aims to increase attention to critical input tokens or decrease attention to less informative tokens. Lu et al. (2021) proposes a mask perturbation method to adjust attention weights for key tokens, thereby improving translation quality. Zhang et al. (2023a) introduces PASTA, which allows manual designation of important tokens during inference. This is extended by AutoPASTA (Zhang et al., 2024a), which uses LLMs to autonomously identify salient tokens and increase attention to them. In contrast, ACT (Yu et al., 2024) reduces attention to semantically trivial sink tokens and redirects it to meaningful content. Similar strategies have been applied to VLMs to mitigate hallucinations. PAI (Liu et al., 2024b) enhances attention to image tokens at inference time to counteract text-dominant bias. IBD (Zhu et al., 2024) and OPERA (Wei & Zhang, 2024) further refine this idea by prioritizing visual information or penalizing overconfident summary tokens. While effective, these methods depend on identifying task-specific tokens, which may introduce bias (e.g., overemphasizing misleading tokens) and limit applicability when token importance is unclear or attention maps are unavailable. In contrast, our method focuses on a task-invariant initial token, removing the need for costly token identification, and can be easily applied by tuning key states.
@@ A.2 The Magic of the Initial Token
Recent studies highlight the significance of the initial token, especially through the lens of the attention sink phenomenon, where it draws substantial attention despite low semantic content. Xiao et al. (2023) shows that preserving such tokens is critical for maintaining performance in sliding window attention. Kaul et al. (2024) attributes this effect to softmax normalization and causal masking, while Gu et al. (2024) and Barbero et al. (2025) identify architectural biases that amplify attention to the initial token, including key-query alignment and LayerNorm effects. Functionally, the initial token is hypothesized to serve as a stabilizing “no-op” anchor, enhancing robustness to prompt variations (Barbero et al., 2025). It has been leveraged in applications such as long-context modeling (Zhang et al., 2023b; Xiao et al., 2023), but also poses challenges for quantization due to its high attention weight (Dettmers et al., 2023; Liu et al., 2024a). While previous work has identified the structural and functional importance of the initial token, its potential as a target for attention tuning remains underexplored. In this work, we provide a detailed analysis of attention tuning of the initial token across layers and heads, demonstrating its consistent influence across different tasks. Our approach bridges the gap between these lines of research by proposing a novel method that advances interpretable attention tuning.
@@ Appendix B Detailed Experimental Setup
Models, Tasks, and Datasets. Models: We evaluate ZeroTuning on four LLMs with distinct attention implementations: Llama-3.1-8B-Instruct (Grattafiori et al., 2024) and Llama-2-13B-Chat (Touvron et al., 2023) with eager attention, Qwen-2-7B (Yang et al., 2024) with SDPA attention, and DeepSeek-R1-14B (DeepSeek-AI et al., 2025) with Flash attention. [^2] Eager, SDPA, and Flash are official attention implementations in modern Transformer libraries. Eager computes the full attention map; SDPA uses PyTorch’s efficient API to select the optimal implementation; Flash relies on fused CUDA kernels from the FlashAttention library. Tasks and Datasets: Our experiments encompass three task types across 15 datasets: (1) Text Classification and Reasoning, including SST-2 (binary sentiment classification) (Socher et al., 2013), SST-5 (fine-grained sentiment analysis) (Socher et al., 2013), MR (movie review polarity detection) (Pang & Lee, 2005), SUBJ (subjectivity classification) (Pang & Lee, 2004), TREC (question type classification) (Li & Roth, 2002), CB (commitment detection) (De Marneffe et al., 2019), and BoolQ (boolean question answering) (Clark et al., 2019); (2) Domain-Specific Multiple-Choice, including MMLU (cross-domain knowledge testing) (Hendrycks et al., 2020), AQUA (math word problems) (Zheng et al., 2024a), MathQA (algebraic reasoning) (Amini et al., 2019), LogiQA (logical reasoning) (Liu et al., 2023), CQA (commonsense reasoning) (Talmor et al., 2018), PIQA (physical commonsense QA) (Bisk et al., 2020), and ARCC (scientific reasoning) (Clark et al., 2018); and (3) Multi-Round Conversation, using MT-Bench (Zheng et al., 2024a).
@@ E.1 Deeper Error Analysis for Unsupervised ZeroTuning
    The Bias Amplification Trap ($\gamma>1$): When increasing the initial token’s attention, the model enters a ”global-integrative” mode. For an initially incorrect, uncertain sample, this often induces a ”competing peaks” phenomenon. Both Figure 12 and 13 illustrate cases where the model first amplifies an existing bias, leading to a deep entropy well corresponding to an incorrect answer. Only with further increases in $\alpha$ does the model’s interpretation ”flip” to the correct one. Figure 13 shows a more complex cascade: after overcoming the format trap, the model first falls into a bias trap (incorrectly predicting ”negative” due to the word ”not”) before finally settling on the correct answer (”positive”). In both cases, if the global entropy minimum is sought, it may lock onto an amplified bias or a formatting error, creating a trap.

### v3

@@ 4.2 Overall Performance of ZeroTuning
For a fair and direct comparison with existing supervised baselines, we focus on supervised ZeroTuning in our main experiments.
@@ 4.3 Unsupervised ZeroTuning
We validate a fully unsupervised variant of ZeroTuning, which removes the need for labeled calibration data by minimizing the model’s average output entropy. Appendix E provides a detailed empirical validation, including visual comparisons, additional analysis, and a breakdown of error patterns. As shown in Figure 7, this entropy-guided method is highly competitive with its supervised counterpart, extending applicability to label-scarce scenarios.
@@ 5 Further Analysis and Key Findings
Finally, we analyze the method’s boundaries and potential side effects. We quantitatively demonstrate that ZeroTuning excels at correcting a model’s uncertain errors but cannot override high-confidence mistakes rooted in flawed pretrained knowledge. We also characterize the negative effects of extreme tuning, which provides a clearer picture of the method’s operational limits (Appendix D). This positions our method as a powerful tool for unlocking a model’s latent knowledge, rather than a substitute for fine-tuning. Intriguingly, we also find that within a safe operational range, the scaling factor can modulate output diversity in a manner analogous to temperature, but with the unique ability to alter the rank-ordering of logits and thereby correct errors that temperature scaling cannot fix. This aligns with our experiments on temperature tuning, where pure temperature adjustment fails to improve benchmark performance, while ZeroTuning yields consistent gains substantially.
6 Further Discussion
6.1 Efficient Calibration and Search
In our current implementation, the offline search for the scaling factor $\gamma$ and head selection with Llama-3.1-8B takes roughly 4–5 minutes on a single GPU, and the inference-time overhead is negligible. We also highlight an interesting empirical observation: the clear U-shaped entropy curve identified in Appendix E suggests that more efficient search strategies, such as Bayesian optimization or simple rule-based early stopping, could further reduce calibration cost. We view this as a promising direction for future optimization.
6.2 ZeroTuning and Supervised Fine-Tuning
Supervised fine-tuning (SFT) remains a strong and versatile way to improve LLMs: by updating parameters, it can simultaneously adjust task knowledge, instruction following, output formatting, and various biases. However, this also means SFT is an indirect optimization over many intertwined behaviors and typically requires more data, compute, and careful hyperparameter tuning.
ZeroTuning takes a complementary approach. Instead of changing parameters, it directly adjusts the attention pattern on the initial token, targeting the specific failure modes identified in our analysis (e.g., format traps and biased attention over context tokens). This makes the intervention more localized and training-free, while still being compatible with SFT.
To study this relationship, we run a controlled experiment on BoolQ where both SFT and supervised ZeroTuning use the same 500 labeled examples. We fine-tune the model with a lightweight LoRA setup (rank $r=4$, 3 epochs) and compare four configurations:
Method BoolQ Accuracy
Vanilla 69.60
Vanilla + SFT (LoRA) 81.20
Vanilla + ZeroTuning (Ours) 82.40
Vanilla + SFT + ZeroTuning (Ours) 83.60
In this limited setting, the training-free ZeroTuning slightly outperforms our small-scale LoRA SFT, and applying ZeroTuning on top of the SFT model yields the best performance. While stronger fine-tuning setups (more data, higher rank, more epochs) may implicitly learn to adjust the initial-token attention in a similar way, we hope our observations provide a useful lens for understanding how SFT optimizes attention patterns and inspire future work on combining training-free and training-based approaches.
@@ 7 Conclusion
This work provides a systematic analysis of how scaling the initial token’s attention affects a Transformer’s attention distribution and downstream accuracy. Building on this analysis, we introduce ZeroTuning, a training-free inference-time method that calibrates a single, task-agnostic token. Across a wide range of tasks and models, ZeroTuning improves performance over vanilla inference and is competitive with (and often stronger than) prior inference-time tuning baselines that require manipulating additional tokens or task-specific heuristics. ZeroTuning supports both supervised calibration (via a small labeled set) and unsupervised calibration (via entropy minimization), and it can be implemented across different optimized attention kernels. Overall, our results highlight the initial token as a simple but effective control point for lightweight, deployment-friendly model adaptation.
@@ A.1 Token-Level Attention Tuning
Token-level attention tuning typically aims to increase attention to critical input tokens or decrease attention to less informative tokens. Lu et al. (2021) proposes a mask perturbation method to adjust attention weights for key tokens, thereby improving translation quality. Zhang et al. (2023a) introduce PASTA, which allows manual designation of important tokens during inference. This is extended by AutoPASTA (Zhang et al., 2024a), which uses LLMs to autonomously identify salient tokens and increase attention to them. In contrast, ACT (Yu et al., 2024) reduces attention to semantically trivial sink tokens and redirects it to meaningful content. Similar strategies have been applied to VLMs to mitigate hallucinations. PAI (Liu et al., 2024b) enhances attention to image tokens at inference time to counteract text-dominant bias. IBD (Zhu et al., 2024) and OPERA (Wei and Zhang, 2024) further refine this idea by prioritizing visual information or penalizing overconfident summary tokens. While effective, these methods depend on identifying task-specific tokens, which may introduce bias (e.g., overemphasizing misleading tokens) and limit applicability when token importance is unclear or attention maps are unavailable. In contrast, our method focuses on a task-invariant initial token, removing the need for costly token identification, and can be easily applied by tuning key states.
@@ A.2 The Magic of the Initial Token
Recent studies highlight the significance of the initial token, especially through the lens of the attention sink phenomenon, where it draws substantial attention despite low semantic content. Xiao et al. (2023) show that preserving such tokens is critical for maintaining performance in sliding window attention. Kaul et al. (2024) attribute this effect to softmax normalization and causal masking, while Gu et al. (2024) and Barbero et al. (2025) identify architectural biases that amplify attention to the initial token, including key-query alignment and LayerNorm effects. Functionally, the initial token is hypothesized to serve as a stabilizing “no-op” anchor, enhancing robustness to prompt variations (Barbero et al., 2025). It has been leveraged in applications such as long-context modeling (Zhang et al., 2023b; Xiao et al., 2023), but also poses challenges for quantization due to its high attention weight (Dettmers et al., 2023; Liu et al., 2024a). While previous work has identified the structural and functional importance of the initial token, its potential as a target for attention tuning remains underexplored. In this work, we provide a detailed analysis of attention tuning of the initial token across layers and heads, demonstrating its consistent influence across different tasks. Our approach bridges the gap between the these lines of research by proposing a novel method that advances interpretable attention tuning.
@@ Appendix B Detailed Experimental Setup
Models, Tasks, and Datasets. Models: We evaluate ZeroTuning on four LLMs with distinct attention implementations: Llama-3.1-8B-Instruct (Grattafiori et al., 2024) and Llama-2-13B-Chat (Touvron et al., 2023) with eager attention, Qwen-2-7B (Yang et al., 2024) with SDPA attention, and DeepSeek-R1-14B (DeepSeek-AI et al., 2025) with Flash attention. [^2] Eager, SDPA, and Flash are official attention implementations in modern Transformer libraries. Eager computes the full attention map; SDPA uses PyTorch’s efficient API to select the optimal implementation; Flash relies on fused CUDA kernels from the FlashAttention library. Tasks and Datasets: Our experiments encompass three task types across 15 datasets: (1) Text Classification and Reasoning, including SST-2 (binary sentiment classification) (Socher et al., 2013), SST-5 (fine-grained sentiment analysis) (Socher et al., 2013), MR (movie review polarity detection) (Pang and Lee, 2005), SUBJ (subjectivity classification) (Pang and Lee, 2004), TREC (question type classification) (Li and Roth, 2002), CB (commitment detection) (De Marneffe et al., 2019), and BoolQ (boolean question answering) (Clark et al., 2019); (2) Domain-Specific Multiple-Choice, including MMLU (cross-domain knowledge testing) (Hendrycks et al., 2020), AQUA (math word problems) (Zheng et al., 2024a), MathQA (algebraic reasoning) (Amini et al., 2019), LogiQA (logical reasoning) (Liu et al., 2023), CQA (commonsense reasoning) (Talmor et al., 2018), PIQA (physical commonsense QA) (Bisk et al., 2020), and ARCC (scientific reasoning) (Clark et al., 2018); and (3) Multi-Round Conversation, using MT-Bench (Zheng et al., 2024a).
@@ E.1 Deeper Error Analysis for Unsupervised ZeroTuning
    The Bias Amplification Trap ($\gamma>1$): When increasing the initial token’s attention, the model enters a ”global-integrative” mode. For an initially incorrect, uncertain sample, this often induces a ”competing peaks” phenomenon. Both Figure 11 and 12 illustrate cases where the model first amplifies an existing bias, leading to a deep entropy well corresponding to an incorrect answer. Only with further increases in $\alpha$ does the model’s interpretation ”flip” to the correct one. Figure 12 shows a more complex cascade: after overcoming the format trap, the model first falls into a bias trap (incorrectly predicting ”negative” due to the word ”not”) before finally settling on the correct answer (”positive”). In both cases, if the global entropy minimum is sought, it may lock onto an amplified bias or a formatting error, creating a trap.
@@
Why Collective Signal Works Better?
The visualizations in Figure 15 reveal a key insight: while individual uncertain samples exhibit highly irregular and sometimes multi-peaked probability/accuracy trajectories as $\gamma$ varies, the aggregate signal constructed from many samples becomes smooth, stable, and strongly aligned with true accuracy improvement.
(a) Uncertain Population ($p<0.5$)
(b) Certain Population ($p>0.5$)
(c) All Samples (Aggregated)
Figure 15: Population-level probability and accuracy trajectories. Gray lines represent individual sample trajectories. Red: average probability of the top token. Blue: average accuracy.
We observe two complementary mechanisms:
  • Statistical Cancellation for $\gamma<1$. Uncertain samples tend to fall into the format trap and display artificially low entropy for invalid outputs, whereas certain samples consistently lose confidence when $\gamma$ decreases. When averaged, these opposing trends cancel out, producing a flat and stable population-level curve. This prevents the collective entropy signal from being dominated by pathological samples.
  • Signal Reinforcement for $\gamma>1$. In this region, both populations—despite their very different behaviors at the individual level—exhibit a shared upward trend in accuracy as $\gamma$ increases. This alignment creates a resonant effect: the aggregated probability and accuracy curves become smoother and strongly monotonic, reliably pointing toward the globally beneficial regime of $\gamma>1$.
Together, these effects explain why Unsupervised ZeroTuning succeeds when optimizing the average entropy. The population average filters out idiosyncratic traps (format errors, bias amplifications) and amplifies the underlying structural signal that reflects the model’s latent reasoning and instruction-following capabilities. This makes the minimum of the average-entropy curve over $\gamma$ a trustworthy indicator of the optimal tuning direction.